Find the exact arc length of the curve


x = e ^ 2t* (sin t + cos t) , y=e^2t(sin t - cos t) ; (−1≤t≤1)

1
Expert's answer
2021-12-06T04:34:38-0500
"x'_t=e^{2t}(\sin t+\cos t+2\cos t-2\sin t)"

"=e^{2t}(3\cos t-\sin t)"


"y'_t=e^{2t}(\sin t-\cos t+2\cos t+2\sin t)"

"=e^{2t}(\cos t+3\sin t)"

"(x'_t)^2+(y'_t)^2=e^{4t}(9\cos^2 t-6\sin t\cos t+\sin ^2 t"

"+\cos^2 t+6\sin t\cos t+9\sin ^2 t)=10e^{4t}"

"L=\displaystyle\int_{-1}^1\sqrt{(x'_t)^2+(y'_t)^2}dt"

"=\displaystyle\int_{-1}^1\sqrt{10e^{4t}}dt"

"=\dfrac{\sqrt{10}}{2}[e^{2t}]\begin{matrix}
 1 \\
 -1
\end{matrix}=\sqrt{10}(\dfrac{e^2-e^{-2}}{2})"

"=\sqrt{10}\sinh(2)\ (units)"